OpenStudy (anonymous):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

OpenStudy (thomas5267):

From the question, we know there is a relationship between the two numbers. Lets call the first one A and the second one B. Try to use mathematical formula to express the following two relationship: 1. The sum of five times a number and twice a second number is 18. 2. The difference of three times the first number and seven times the second number is -22.

OpenStudy (anonymous):

5x + 2y = 18 3x - 7y = -22 Is that correct @thomas5267

OpenStudy (thomas5267):

Correct. Do you know how to use the elimination method?

OpenStudy (anonymous):

Uhh not really @thomas5267

OpenStudy (thomas5267):

For example: \[ \begin{align*} 4x + 5y &= 14 \\ -4x - 3y &= -10 \end{align*} \] The elimination method is when you add or subtract one equation to another equation to eliminate one variable, like this. \[ \begin{align*} (4x+5y)+(-4x-3y)&=(14)+(-10) \\ 2y&=4 \\ y&=2 \end{align*} \]

OpenStudy (anonymous):

Ohhhh ok

OpenStudy (anonymous):

ugh im still kinda confused @thomas5267

OpenStudy (thomas5267):

Hint: \(5x+2y=18\) and \(15x+6y=54\) is the same.

OpenStudy (anonymous):

So is the answer C?

OpenStudy (anonymous):

@thomas5267

OpenStudy (anonymous):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. Let the number be x and second number be y. So: 5x + 2y = 18 But 3x - 7y = -22 Now you have two equations. To solve by elimination method, you need to bring one of the coefficients (either of x or y) equal for both equations then cancel. 5x + 2y = 18 Multiply everything by 3 3x - 7y = -22 Multiply everything by 5 15x + 6y = 54 15x - 35y = -110 Subtract 15x will get cancelled. So, 41y = 164 y = 164/41 = 4 So, y = 4 Substitute this value of y=4 in equation 1. 5x + 2(4) =18 5x + 8 = 18 5x = 18 - 8 5x = 10 x = 10/5 =2 So, x = 2 and y = 4. That is D.